# k-LLMmeans: Summaries as Centroids for Interpretable and Scalable LLM-Based Text Clustering

### v1

@@ 6 Experiments
Table 1: Average Normalized Mutual Information (NMI) and distance between final and true centroids (dist) for k-means, k-medoids, and four k-LLMmeans variants across 10 random seeds on four benchmark datasets (including both domain and intent from MASSIVE), using four different embedding models.
6.1 Datasets
We evaluate our clustering approach on four benchmark datasets:
  • Bank77 (Casanueva et al., 2020): Consists of 3,080 customer queries related to banking services, categorized into 77 distinct intents.
  • CLINC (Larson et al., 2019): A diverse set of 4,500 queries spanning 150 intent classes across multiple domains, designed for open-domain intent classification.
  • GoEmo (Demszky et al., 2020): Contains 2,984 social media posts annotated with 27 fine-grained emotion categories. We removed the neutral expressions to address data imbalance and retained only entries with a single, unique emotion.
  • MASSIVE (FitzGerald et al., 2023): Comprises 2,974 English-language virtual assistant utterances grouped into 18 domains and 59 intent categories.
6.2 New compiled dataset for testing text-Streaming Clustering Algorithms
We extract and unify a challenging data stream comprising unique archive posts collected from 84 Stack Exchange sites (StackExchange, 2024). Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods. Our raw dataset spans 84 domains, each containing at least 20 posts per year from 2018 to 2023 (with post lengths ranging from 20 to 1000 characters), totaling 499,359 posts. For our experiments, we focus on posts from 2020 to 2023 and further filter out labels that do not exceed 500 posts in 2023. The resulting subset comprises 35 distinct groups and 69,147 posts. Both the raw and clean data are provided with this paper.
6.3 Methods
@@
NMI dist NMI dist NMI dist NMI dist
k-means 80.6 0.14 79.0 0.167 79.0 0.154 79.6 0.138
mini-batch k-means 78.2 0.181 77.4 0.166 77.6 0.173 77.0 0.175
seq. mini-batch k-means 76.6 0.187 75.2 0.185 75.6 0.187 74.8 0.184
mini-batch k-LLMmeans-1 81.2 0.141 80.3 0.138 79.4 0.145 80.0 0.136
mini-batch k-LLMmeans-5 80.9 0.144 80.5 0.127 80.5 0.125 80.3 0.129
mini-batch k-LLMmeans-FS1 81.1 0.136 79.8 0.141 79.3 0.147 79.8 0.138
mini-batch k-LLMmeans-FS5 81.6 0.126 80.2 0.126 80.1 0.129 80.1 0.133
Table 2: Average Normalized Mutual Information (NMI) and distance from final and true centroids (dist) for k-means, mini-batch k-means, sequential mini-batch k-means and four sequential mini-batch k-LLMmeans variants with across 5 random seeds on StackExchange data.
We evaluate our k-LLMmeans algorithm on each of the four static (non-streaming) dataset using the known number of clusters and performing 120 centroid-update iterations and 10 different seeds. To demonstrate the robustness of our approach, we compute embeddings with four different pretrained models: distilbert (Sanh, 2019), e5-large (Wang et al., 2022), s-bert (Reimers and Gurevych, 2019), and OpenAI’s text-embedding-3-small (OpenAI, 2023). For the LLM component, we only use OpenAI’s gpt-4o to ensure any observed differences in performance arise from the effectiveness of our clustering method rather than the inherent strengths or weaknesses of different LLMs.
For the instruction task $I$, we employ a simple summarization prompt that adapted to each dataset. For example, for Bank77 we use the prompt: “The following is a cluster of online banking questions. Write a single question that represents the cluster concisely.” We examine four variations of k-LLMmeans based on different numbers of summarization steps and the size of the prompts:
  • k-LLMmeans-1: A single summarization step ($l=60$) using all documents in the cluster as input.
  • k-LLMmeans-5: Five summarization steps ($l=20$) with all documents in the cluster as input at each step.
  • k-LLMmeans-FS1: A few-shot variant with 10 randomly selected documents in a single summarization step.
  • k-LLMmeans-FS5: A few-shot variant with 10 randomly selected documents in five summarization steps.
Our baselines are the standard centroid-based algorithms k-means and k-medoids. While alternative clustering methods may achieve stronger performance, our primary goal is to demonstrate improvement specifically over widely used centroid-based approaches.
Methods for streaming dataset. We partition the StackExchange data into four subsets, each corresponding to a single year from 2020 to 2023 (See Table 2 for yearly dataset sizes). For each yearly subset, we split the data into $b=\lceil\tfrac{n}{10000}\rceil$ equal-sized batches ${D_{1},\dots,D_{b}}$ in chronological order, where $n$ is the number of documents for that year. We then use the ground-truth clusters and run the mini-batch k-LLMmeans algorithm in its four previously described variants (for practical reasons for the full cluster method we use 50 samples). For comparison, we evaluate three baselines: mini-batch k-means with standard random sampling across the entire year (until convergence), sequential mini-batch k-means with $b$ chronological batches, and standard k-means on the full dataset. In all experiments, we use OpenAI’s models for both embeddings and the LLM.
6.4 Results
Table 1 reports the average Normalized Mutual Information (NMI) across all method–dataset–embedding combinations, as well as the average distance (dist) between the centroids produced by each method and the true centroids calculated with the ground truth clusters. A lower distance indicates greater proximity to the optimal solution. This distance aligns well with the objectives of our approach. We also measure Accuracy (not shown here due to space constraints), observing results consistent with the reported metrics. Overall, all k-LLMmeans variants outperform k-means except when using the distilbert embedding. However, the performance of every method using this embedding is clearly inferior compared to those employing any other embedding. Meanwhile, k-medoids emerges as the worst performer. Comparing the four k-LLMmeans configurations reveals only minor differences, although running five summarization steps ($l=20$) tends to offer better performance. Interestingly, using few-shot summarization generally leads to better results than prompting on the entire cluster. These observations suggest k-LLMmeans improves over k-means while remaining efficient, as few-shot summarization appears sufficient without needing extensive LLM usage.
@@ 6.4 Results
We present the same evaluation metrics for sequential data in Table 2. The mini-batch k-LLMmeans methods consistently outperform all three baselines across all scenarios. Notably, they even surpass standard k-means, which operates on the full dataset rather than mini-batches. This highlights the scalability and efficiency of our approach for sequential data, achieving superior performance despite processing data sequentially. As observed previously, few-shot summarization proves effective, mitigating the need for extensive LLM usage.
@@ 7 Case study
Figure 2: Sequential evolution of the LLM-generated centroids for four primary clusters during the three batches of the sequential mini-batch k-LLMmeans process applied to 2021 posts from the AI Stack Exchange site (StackExchange, 2024). Main aspects are manually highlighted in different color on each cluster.
To demonstrate the interpretability of our method in capturing the evolution of clusters within sequential data, we present a case study using posts from the AI site in the 2021 Stack Exchange dataset (StackExchange, 2024). We apply our mini-batch k-LLMmeans algorithm with three equal-length batches and a total of ten clusters. We use the instruction “The following is a cluster of questions from the AI community. Write a single question that represents the cluster”.
@@
8 Conclusions
@@
We introduced k-LLMmeans and mini-batch k-LLMmeans, unsupervised clustering algorithms that refine k-means by integrating a lightweight LLM-based summarization into centroid updates. This modification enriches the contextual representation while preserving k-means’ efficiency and scalability, yielding interpretable clusters for both static corpus of documents and streaming text data.
Limitations

### v2

@@ 5.1 Results with static data
Table 2: Average ACC, NMI, and distance between final and true centroids ($dist$) for k-means and k-LLMmeans FS-multiple, evaluated on three datasets using four different embedding models.
Dataset CLINC GoEmo Massive (D)
embedding/Method ACC NMI $dist$ ACC NMI $dist$ ACC NMI $dist$
DistilBERT
   k-means 53.9 77.2 0.340 17.8 18.3 0.364 44.4 45.1 0.309
   k-LLMmeans 55.3 78.7 0.343 18.2 18.8 0.351 46.2 46.4 0.295
text-embedding-3-small
   k-means 77.3 92.0 0.200 20.7 20.5 0.287 59.4 67.9 0.246
   k–LLMmeans 80.2 93.1 0.179 24.2 22.3 0.278 63.2 70.6 0.227
e5-large
   k-means 73.8 90.8 0.131 22.8 22.8 0.176 58.4 63.7 0.138
   k-LLMmeans 77.2 92.5 0.119 24.2 24.3 0.168 62.3 65.9 0.133
S-BERT
   k-means 76.9 91.0 0.215 13.7 13.3 0.355 58.2 64.6 0.271
   k-LLMmeans 79.7 92.5 0.198 14.7 13.9 0.346 59.7 65.6 0.254
Table 3: Number of LLM calls (prompts), average ACC, and average NMI for k-LLMmeans FS-multiple using various LLMs and e5-large embeddings, compared against other LLM-based text clustering methods on three benchmark datasets.
Dataset/Method CLINC GoEmo Massive (D)
prompts ACC NMI prompts ACC NMI prompts ACC NMI
k-LLMmeans
   GPT-3.5 750 76.0 92.1 135 23.9 24.1 90 63.3 66.2
   GPT-4o 750 77.2 92.5 135 24.2 24.3 90 62.3 65.9
   Llama-3.3 750 77.2 92.3 135 23.8 23.1 90 62.0 66.3
   DeepSeek-V3 750 69.7 90.8 135 22.8 22.7 90 62.6 66.3
   Claude-3.7 750 76.9 92.5 135 24.2 23.7 90 61.8 66.0
ClusterLLM 1618 83.8 94.0 1618 26.8 23.9 1618 60.9 68.8
IDAS 4650 81.4 92.4 3011 30.6 25.6 2992 53.5 63.9
LLMEdgeRefine 1350 86.8 94.9 895 34.8 29.7 892 63.1 68.7
Comparing k-LLMmeans variants and traditional clustering methods. Table 1 reports Average Accuracy (ACC) and Normalized Mutual Information (NMI) for four k-LLMmeans variants using GPT-4o, compared against traditional clustering baselines, with text-embedding-3-small embeddings, evaluated on four benchmark datasets. Overall, all k-LLMmeans variants outperform traditional clustering methods in terms of NMI and in general achieve higher average accuracy across datasets. Within k-LLMmeans variants there are minor differences, although running multiple summarization steps tends to offer better performance. Interestingly, using few-shot summarization generally leads to better results than prompting on the entire cluster. These observations suggest k-LLMmeans improves over k-means while remaining efficient, as few-shot summarization appears sufficient without needing extensive prompt length (See Appendix B for standard errors of all simulations).
[…]
Comparing k-LLMmeans with different LLMs and state-of-the-art LLM-based clustering methods. Table 2 reports number of LLM calls (prompts) and Average Accuracy (ACC) and Normalized Mutual Information (NMI) for the k-LLMmeans FS-multiple variant on three benchmark datasets, pairing e5-large embeddings with several backbone LLMs. Performance is stable across LLM choices, indicating that k-LLMmeans is robust to the specific language model. For reference, Table 2 also reports results from Feng et al. [15] for ClusterLLM, IDAS, and LLMEdgeRefine. All three baselines rely on GPT-3.5: ClusterLLM (with fine-tuning) and IDAS use the same e5-large embeddings as we do, whereas LLMEdgeRefine benefits from the stronger Instructor embeddings [21]. Even so, our GPT-3.5 configuration attains ACC and NMI that are only marginally lower—and often comparable—while requiring orders of magnitude fewer LLM calls and no fine-tuning. This substantial reduction in model prompts translates directly into improved scalability and cost-efficiency.
6 Sequential (mini-batch) experiments
We use yearly posts from 2020 to 2023 from 35 StackExchange sites [41]. Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods (See Appendix A.2 for details). For each yearly subset, we split the data into $b=\lceil\tfrac{n}{10000}\rceil$ equal-sized batches ${D_{1},\dots,D_{b}}$ in chronological order, where $n$ is the number of documents for that year. We then use the ground-truth clusters and run the mini-batch k-LLMmeans algorithm in its four variants described in Section 5 (for practical reasons for the full cluster method we use 50 samples). For comparison, we evaluate three baselines: mini-batch k-means with standard random sampling across the entire year (until convergence), sequential mini-batch k-means with $b$ chronological batches, and standard k-means on the full dataset. In all experiments, we use text-embedding-3-small and GPT-4o for the summarization step.
6.1 Results with streaming data
Table 4: Average ACC, and average NMI for four sequential mini-batch variants, k-means, mini-batch k-means, sequential mini-batch k-means on the yearly StackExchange data.
@@ 6.1 Results with streaming data
ACC NMI ACC NMI ACC NMI ACC NMI
mini-batch k-LLMmeans
   single 73.1 81.2 73.2 80.3 70.8 79.4 72.3 80.0
   multiple 72.5 80.9 73.6 80.5 74.4 80.5 73.4 80.3
   FS + single 74.0 81.1 72.2 79.8 70.2 79.3 71.6 79.8
   FS + multiple 75.4 81.6 73.5 80.2 72.8 80.1 72.7 80.1
k-means 73.4 80.6 67.7 79.0 68.6 79.0 72.0 79.6
mini-batch k-means 67.0 78.2 67.7 77.4 67.5 77.6 67.2 77.0
seq. mini-batch k-means 67.0 76.6 66.7 75.2 65.6 75.6 65.8 74.8
@@
Table 4 presents ACC and NMI on the yearly StackExchange corpus, comparing four sequential mini-batch variants of k-LLMmeans against standard k-means, mini-batch k-means, and sequential mini-batch k-means across five random seeds. The mini-batch k-LLMmeans methods consistently outperform all three baselines across all scenarios. Notably, they even surpass standard k-means, which operates on the full dataset rather than mini-batches. This highlights the scalability and efficiency of our approach for sequential data, achieving superior performance despite processing data sequentially. As observed previously, few-shot summarization proves effective, mitigating the need for extensive LLM usage.
@@ 7 Case study
To demonstrate the interpretability of our method in capturing the evolution of clusters within sequential data, we present a case study using posts from the AI site in the 2021 Stack Exchange dataset [41]. We apply our mini-batch k-LLMmeans algorithm with three equal-length batches and a total of ten clusters. We use the instruction “The following is a cluster of questions from the AI community. Write a single question that represents the cluster”.
@@
Figure 3: Sequential evolution of the LLM-generated centroids for four primary clusters during the three batches of the sequential mini-batch k-LLMmeans process applied to 2021 posts from the AI Stack Exchange site [41]. Main aspects are manually highlighted in different color on each cluster.
8 Related work
Traditional clustering. Hierarchical methods [22, 5] build tree-structured representations of nested document relationships. Density-based approaches like DBSCAN [14] and graph-based methods detect clusters of arbitrary shapes, while spectral clustering [29] leverages eigen-decomposition to uncover complex structures. Model-based techniques—including Gaussian mixture models [11] and recent neural network frameworks [52, 18, 49, 50, 48]—provide probabilistic clustering formulations. Additionally, topic modeling methods, from probabilistic latent semantic analysis [17] to latent Dirichlet allocation [6], capture word co-occurrence patterns and latent topics.
LLM-based clustering. Viswanathan et al. [45] employ LLMs to augment document representations, generate pseudo pairwise constraints, and post-correct low-confidence assignments for query-efficient, few-shot semi-supervised clustering. Zhang et al. [51] propose ClusterLLM, which uses instruction-tuned LLMs via interactive triplet and pairwise feedback to cost-effectively refine clustering granularity. Complementary approaches [44, 33] show that context-derived representations capture subtle semantic nuances beyond traditional embeddings. Wang et al. [47] introduce a goal-driven, explainable method that employs natural language descriptions to clarify cluster boundaries, while [10] present IDAS for intent discovery using abstractive summarization. Feng et al. [15] propose LLMEdgeRefine an iterative mechanism that forms super-points to mitigate outliers and reassign ambiguous edge points, resulting in clusters with higher coherence and robustness.
9 Discussion
@@
We introduced k-LLMmeans and mini-batch k-LLMmeans, unsupervised text clustering algorithms that refine k-means by integrating a lightweight LLM-based summarization into centroid updates. This modification enriches the contextual representation while preserving k-means’ stability and scalability, yielding interpretable clusters for both static corpus of documents and streaming text data, with low LLM usage.